OpenStudy (anonymous):

Divide. (-6z^7u+7z^7u^6)÷(-2z^5u^3) Simplify your answer as much as possible.

OpenStudy (anonymous):

\[(-6z ^{7}u+7z ^{7}u ^{6})\div(-2z ^{5}u ^{3})\] Why does mathway and symbolab both have different answers.

OpenStudy (anonymous):

-1/2z^9u^4

OpenStudy (anonymous):

If u don't understand it I will not mind to explain to

OpenStudy (anonymous):

U how I got the answer

OpenStudy (anonymous):

Yes can you explain it for me. Thank you.

OpenStudy (anonymous):

using the laws of indices

OpenStudy (anonymous):

a*a=a² a+a=2a

OpenStudy (anonymous):

in division its the reverse

OpenStudy (anonymous):

so all u have to do is apply the law

OpenStudy (anonymous):

-6z⁷u+7z⁷u⁶ -6+7=-1 z⁷+z⁷=z¹⁴ u+u³=u⁴ =-1z¹⁴u⁴

OpenStudy (anonymous):

-6+7 = 1 ? or -1

OpenStudy (anonymous):

Thank you 1

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

normal answer is 1z^14u^7/2z^5u^3

OpenStudy (anonymous):

Now u can simplify that

OpenStudy (anonymous):

=1/2z^9u^4

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }u ^{10}z ^{19}\]